You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100218 (09 Apr 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 February17-18

Observer: V. Testa, K.Boutsia, F. Pedichini
Telescope Operator: D.Gonzales Huerta
Support Astronomer: D. Thompson
Instrument: LUCIFER1 + LBC

Summary

Started with Lucifer and lost ~1hr for mirror collimation and other similar problems
Switched to LBC, It took only 20 min
Made some tests to close AO loop online with A. Baruffolo in remote. Continuing next night

Data Summary

Project Objects Filter ExpTime Notes
Longhetti MUNICS 150_Ks 1 hr Long Slit, telluric taken
Fiore Cosmos Uspec, g, r 4 hr Taken g,r in binocular
Mereghetti LDRT Uspec 0.5 hr  

Details

01:16 - Pointing Star High Elevation (WT10_314) 0002

01:21 - Pointing Star Low Elevation (WT10_078) 0003

01:26 - Point Star High Elevation (WT10_314) 0004

01:30 - Persson standard BS9105 for collimation (POSITION,ACTIVE)

01:49 - Collimation - preset to Longhetti ACQSCI script

01:50 - Hit primary mirror limits - redo aquisition

01:53 - redo preset and aquisition

01:59 - seems to be working - AQCSCI script
  • S2F1-142 - Ks 3x10x12 with jitter 0005 - 0009

02:02 - Exposure aborted for a PSF kind of problem - guide star out of the field ????

02:14 - Pointing star again for pointing correction ACT0210

02:18 - Starting Longhetti SCI script and begin with the spectroscopy * S2F1-142 - Ks 300x1x12 with jitter 0011 - 0026

03:38 - Flat acquisition script (Calibration Unit on) 0027-0042

03:45 - Arc acquisition script ??? nothing written on header!!!!
  • Dark Ar 0047
  • ARC Ar 0048
  • Dark Ne 0049
  • ARC ne 0050

03:50 Correct Pointing 0051

03:53 - Slew to Hip 24311 - (TEL2 ACQ script) 0052-0055
  • Telluric Offset position acquisition aborted due to rotator failure 20x3 integrated
  • To resolve problem telescope needs to be brought to Zenith and reset manually (aprox. time needed 30 minutes)

04:08 - We decide to switch from LUCIFER to LBC

04:39 - Start using LBC - slew to pointing star ACT0231 - dofpia, /Blueonly
  • We apply Z7= -3000 and Z8 =-1000
  • -13366 0 0 0 0 0 0 0.61 5.27
  • 958 -89 -861 90 -528 -360 48 1.03 0.57
  • 218 -168 -449 -94 -116 -141 45 1.12 0.21
  • Pointing correction CA, IE : -22.9, -31.1

04:47 - Slew to COSMOS - dofpia, /Blueonly

04:56 Collimation not convincing - dofpia, /Blue
  • LBC Controller Error

05:05 - dofpia, /Blueonly
  • -210 453 299 79 74 142 0 0.87 0.22
  • -99 192 349 -12 79 48 -16 0.89 0.15

05:08 - cosmos_Fiore_Uspec_P2_c

05:44 - dofpia, /Blue
  • NO GOOD PUPILS FOUND
  • dofpia, /Blue, /X2

05:47 - Shutter delay- Shutter remained closed and it had to be manually reset

06:03 - Shutter problem fixed - dofpia, /Blue
  • -966 90 570 48 90 253 -48 0.83 0.46
  • -12 189 -326 -65 -13 46 -59 0.84 0.14

06:06 - cosmos_Fiore_Uspec_P2_d

06:41 - dofpia, /Blue
  • 60 -503 338 27 -46 -33 -23 0.95 0.20
  • 822 340 -73 -108 -17 -174 13 1.30 0.37
  • 347 167 -430 37 2 -53 27 1.44 0.21

06:46 - cosmos_Fiore_Uspec_P2_e

07:08 - Shutter blocked - (cosmos_Fiore_Uspec_P2_e stopped)

07:22 - Worked around but probably unstable, we suspect mechanical problems
  • Recollimated and restarted the OB

07:34 - Again shutter problems - OB stopped

Worked on it restarting lbckill/lbcstart then trying with rotating at PA=180

07:47 - Start Cosmos_Fiore_Uspec_P2_e_180
  • A bit out of focus, correcting by hand: Z4=-500, then again -500 and -200
  • first two images seem to be ok
  • At 180 images fine

08:20 - dofpia,/BlueOnly
  • 775 -117 -63 64 45 397 28 0.56 0.41
  • -71 -248 -508 47 14 66 15 0.78 0.19
  • -233 -59 273 -211 76 134 3 0.69 0.20

08:26 - cosmos_Fiore_Uspec_P2_f_180

09:03 - dofpia,/BlueOnly
  • -796 -92 -60 131 -375 41 23 0.65 0.37
  • -258 -217 106 2 -55 -24 -10 0.67 0.15

09:07 - cosmos_Fiore_Uspec_P3_a

09:40 - dofpia,/BlueOnly -175 -230 92 73 59 -12 28 0.76 0.13

09:42 - cosmos_Fiore_Uspec_P3_b

09:18 - Started using both mirrors -dofpia

R: 3284 -82 81 -269 113 434 112 0.55 1.33
B: -277 247 -350 65 258 212 6 0.84 0.26

R: -3953 153 130 40 -47 855 132 0.58 1.65
B: -258 -163 147 55 4 31 17 0.86 0.15

R: -1923 -18 51 -39 -10 352 87 0.47 0.79
B: -265 200 -347 28 37 37 21 0.81 0.18

R: -1030 56 -102 34 77 264 56 0.52 0.45
B: -302 -69 127 -25 9 78 17 0.84 0.16

R: -789 44 102 156 -30 199 34 0.55 0.35
B: -227 310 -190 48 -48 25 -10 0.92 0.16

R: -342 136 116 -121 3 109 32 0.60 0.19
B: 182 -254 105 -53 -7 -12 3 0.95 0.14

10:29 - cosmos_Fiore_gr_P1_a

11:10 Technical Session - Adaptive Optics Test - Slew to M5 - dofpia

R 1699 -169 -203 -36 -160 -183 -81 0.59 0.69
B 2241 442 414 -194 134 -40 -39 0.92 0.91

R 540 3 114 65 -109 -140 -32 0.56 0.25
B 521 248 -64 19 21 -48 -3 0.87 0.23

R 303 -57 -116 -105 60 -99 -20 0.50 0.17
B 17 238 75 -126 -25 -35 -31 0.78 0.12

11:17 - AOTests_M5 *dofpia

R 597 113 -12 -15 -158 -193 -48 y 0.48 0.29
B 513 -331 58 -68 -84 -17 -3 y 0.86 0.24

R 284 -241 -59 76 37 -20 11 y 0.52 0.16
B 134 258 -92 -70 -25 0 -27 y 0.80 0.13

12:23 - Slew to LDRT (Mereghetti) - dofpia

R -719 182 59 292 -22 -119 -3 0.42 0.34
B -774 -220 -314 333 -202 -55 -13 0.83 0.38

R 179 165 -89 -43 -54 -90 26 0.42 0.14
B 291 -169 -24 -95 -29 -180 9 0.79 0.19

12:29 - LDRT-Ui-3La
  • Z4=+500 - red channel
  • second dofpia after first exposure (OB canceled)

R 3536 221 -25 -63 -145 651 97 0.54 1.46
B 89 -11 -157 32 -64 135 -6 0.88 0.13

R -340 162 23 58 -29 127 62 0.57 0.18
B -167 197 64 1 115 70 -13 0.89 0.14

12:43 - LDRT-Ui 3La (OB resumed)

13:16 - SkyFlatTest_gr

13:19 - SkyFlat_gr_5 : 5sec

13:23 - SkyFlat_gr_180_5 : 3sec

13:29 - SkyFlatTest_UsY

13:31 - SkyFlat_UsY_5 :3sec

13:35 - SkyFlat_UsY_5_rot180 :2sec

13:40 - LUCIFER DARK
  • 1x300sx10
  • 1x30sx10

13:50 - End of Night

-- VincenzoTest - 17 Feb 2010
Topic revision: r5 - 09 Apr 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback